OpenStudy (rockycatsrule):

What is the slope of the line passing through the points (–1, 3) and (4, –7)? A. 2 B. 3/4 C. -4/3 D. -2 i dont get this one

rvc (rvc):

\[m=\frac{ y_2-y_1 }{ x_2-x_1 }\]

rvc (rvc):

(-1,3)=(x1,y1) (4,-7)=(x2,y2)

OpenStudy (liv1234):

Do you k now how to solve for at the slope?

OpenStudy (liv1234):

\[\frac{ -1-3 }{ 4-7 }=\frac{ -4 }{ }\]

OpenStudy (liv1234):

There's supposed to be a negative 3 under that.

OpenStudy (liv1234):